OpenStudy (anonymous):

Using the quadratic formula, identify the a, b and c values from the equation: x^2=4x+1=0

OpenStudy (goten77):

x^2=4x+1=0 um should it be x^2-4x+1=0 ??? or x^2=4x+1 then set that equal to 0 and solve for x? *thers 2 equal signs*